# Fundamental limitations on the recoverability of quantum processes

### v1

@@ IV Sufficiency for the recoverability of quantum channels
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}^{2}_{1}$ is the fidelity of states $\rho$ and $\sigma$, $\mathsf{C}_{\mathcal{N}}^{\Psi}$ is the Choi state for channel $\mathcal{N}_{A\rightarrow A^{\prime}}$ with respect to $\Psi$ and $\mathcal{P}^{\mathrm{R}}$ is the recovery map defined by Eq. (70). $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{D}(C\otimes C^{\prime})$ is some fixed state and $\mathfrak{T}:\mathcal{L}\left(A\otimes A^{\prime}\right)\rightarrow\mathcal{L}\left(C\otimes C^{\prime}\right)$ is the representing map for superchannel $\Theta$ (see Remark 3 and Eq. (29)).
The detailed proof of this theorem is provided in Appendix  D while we sketch the proof of this theorem in the following. The basic idea behind the proof is the fact that a supermap induces a linear map, called the representing map, at the level of Choi matrices. The representing map $\mathfrak{T}$ of the superchannel $\Theta$ is in general a CP map that becomes trace-preserving if the adjoint of $\Theta$ satisfies a simple condition (see Eq. (109) of Appendix III). This, in trun, implies that $\Theta^{*}$ is a completely depolarising map preserving (see Lemma 6) if $\Psi$ and $\Phi$ are maximally entangled states. Even if this trace-preserving condition of the representing map $\mathfrak{T}$ is not satisfied, we can always replace the representing map $\mathfrak{T}$ with the following CPTP map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{D}(C\otimes C^{\prime})$ is some fixed state and for all $\mathsf{X}\in\mathcal{L}(A\otimes A^{\prime})$, such that we have $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$ and $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{M}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{M}})$. The recovery map $\mathcal{P}^{\mathrm{R}}$ is defined in such a way that it always recovers $\mathsf{C}^{\Psi}_{\mathcal{M}}$ from $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{M}})$.
@@
[…]
@@
Let $\mathcal{N},\mathcal{M}\in\mathcal{L}\left(A,A^{\prime}\right)$ be two tele-covariant quantum channels (with respect to the same unitary representation of a group) (DW19a, , Definition 4), and let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a tele-covariance preserving superchannel. Then we have
@@
Let $\mathcal{N}\in\mathcal{L}\left(A,A^{\prime}\right)$ be a tele-covariant channel and $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a tele-covariant channel preserving superchannel. Then change in the entropy of quantum channel $\mathcal{N}$ is bounded as
@@
where $\mathsf{C}_{\mathcal{N}}$ is the Choi state for channel $\mathcal{N}$ and $\mathsf{T}\in\mathcal{L}\left(A\otimes A^{\prime},C\otimes C^{\prime}\right)$ is the representing map for superchannel $\Theta$ (see Remark 3 and Eq. (29)). $\mathsf{T}^{\prime}(\mathsf{X}):=\mathsf{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{D}(C\otimes C^{\prime})$ is some fixed state. The recovery map $\widetilde{\mathcal{P}}^{\mathrm{R}}\left(\mathsf{X}\right)$ (see (BDW, 16, Eq. (36))) is given by $\widetilde{\mathcal{P}}^{\mathrm{R}}\left(\mathsf{X}\right)={\mathsf{T}^{\prime}}^{*}(\mathsf{X})+\operatorname{tr}\left(\left(\mathrm{id}-{\mathsf{T}^{\prime}}^{*}\right)\mathsf{X}\right)\xi$, where $\xi\in\mathcal{D}(C\otimes C^{\prime})$, $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$.
@@
Since $\mathcal{N}$ is a tele-covariant channel and $\Theta$ is a tele-covariant channel preserving superchannel, we have $S\left[\mathcal{N}\right]=S(\mathsf{C}_{\mathcal{N}})-\log\left(|A|\right)$ and $S\left[\Theta\left(\mathcal{N}\right)\right]=S\left(\mathsf{C}_{\Theta(\mathcal{N})}\right)-\log\left(|C|\right)$, where $\mathsf{C}_{\mathcal{N}}$ and $\mathsf{C}_{\Theta(\mathcal{N})}$ are the Choi states of $\mathcal{N}$ and $\Theta(\mathcal{N})$, respectively. Then the entropy difference becomes $S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]=S\left(\mathsf{C}_{\Theta(\mathcal{N})}\right)-S(\mathsf{C}_{\mathcal{N}})+\log\widetilde{d}$, where $\widetilde{d}:=|A|/|C|$. Also, $\mathsf{C}_{\mathcal{N}}$ and $\mathsf{C}_{\Theta(\mathcal{N})}$ are connected by the representing map $\mathsf{T}$ which is CP but not trace preserving, in general. So we replace it with a channel $\mathsf{T}^{\prime}$ such that $\mathsf{T}^{{}^{\prime}}(\mathsf{C}_{\Psi})=\mathsf{T}(\mathsf{C}_{\Psi})$ holds when $|A|=|C|$ (see Eq. (20)). The entropy difference then is lower bounded by (see (BDW, 16, Eq. (37)), also DKSW (18))
@@ V Generalized divergence and entropy of superchannels
In Section III, we generalized the definition of entropy to quantum channels, where we argued that similar to the entropy of quantum states, the entropy of quantum channels can also be defined in terms of relative entropy between the given quantum channel and the completely depolarising map (see Definition 46). Here, we first extend the definition of generalized divergences to superchannels using the generalized divergence between two quantum channels. Recall that a map $\mathbf{D}:\mathcal{L}\left(A\right)_{+}\times\mathcal{L}\left(A\right)_{+}\rightarrow\mathbb{R}$ is called a generalized state divergence if it satisfies data processing inequality: $\mathbf{D}\left(\rho\;\middle\|\;\sigma\right)\geq\mathbf{D}\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)$, where $\rho,\sigma\in\mathcal{L}\left(A\right)_{+}$ and $\mathcal{N}:\mathcal{L}(A,A)$ is a quantum channel. Given a generalized state divergence $\mathbf{D}$, we define the generalized divergence between a quantum channel $\mathcal{N}\in\mathcal{L}\left(A,A^{\prime}\right)$ and a CP map $\mathcal{M}\in\mathcal{L}\left(A,A^{\prime}\right)$ as
@@
where $R$ is a reference system of arbitrary dimension. Now, given a generalized divergence $\mathbf{D}$ between channels (cf. CMW (16); LKDW (18); Yua (19)), we define the generalized divergence between two quantum superchannels $\Theta$ and $\Gamma$, denoted by $\mathbf{D}^{(2)}$, as follows.
@@
Given a generalized divergence $\mathbf{D}\left[\cdot\;\middle\|\;\cdot\right]$ between a quantum channel and a CP map (as in Eq. (77)), the generalized divergence between a quantum superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ and a completely CP-preserving supermap $\Gamma:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is defined as
@@
where $\Lambda_{\text{pre}}\in\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(A\otimes R,B\otimes R\right)\right]$ and $\Lambda_{\text{post}}\in\mathscr{L}\left[\mathcal{L}\left(C\otimes R,D\otimes R\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ are preprocessing and postprocessing superchannels, respectively, similar to the preprocessing and postprocessing quantum channels of Eq. (13). $R$ is the reference system on which $\operatorname{id}$ acts. Looking at the mathematical similarity to quantum channels, we expect that the dilation equality, Eq. (79), will hold for all quantum super-superchannels. But, for the rest of this section, we restrict ourselves to the class $\mathsf{SSChan}$ of super-superchannels. For any pair of superchannels $\Theta,\Gamma\in\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]$, our definition of the generalized divergence $\mathbf{D}^{(2)}$ for superchannels satisfies the monotonicity relation $\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\Gamma\right]\geq\mathbf{D}^{(2)}\left[\mathring{\mathbf{\Lambda}}\left(\Theta\right)\;\middle\|\;\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right]$ for all $\mathring{\mathbf{\Lambda}}\in\mathsf{SSChan}$. The proof goes as follows. For $\mathring{\mathbf{\Lambda}}\in\mathsf{SSChan}$, the superchannels $\mathring{\mathbf{\Lambda}}(\Theta)$ and $\mathring{\mathbf{\Lambda}}(\Gamma)$ can be written in terms of concatenation of superchannels given by Eq. (79). Using our definition of divergence between two superchannels, we can write
@@
In the following, we define a notion of generalized entropy for quantum superchannels based on the notion of the generalized divergence. Recall that generalized divergences $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ are functionals that take a pair of operators, say $A,B\geq 0$, as input and give a real number as output. One can define generalized entropies by taking $B$ to be the identity operator. For example, by taking $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ as relative entropy, and $B=\mathbbm{1}$, we get von Neumann entropy, which we have already discussed in the previous section. Also, by taking $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ as sandwiched Renyi divergence WWY (14); ${}^{+}$ (13), and again $B=\mathbbm{1}$, one can get $\alpha$-Renyi entropy. To define the entropy of quantum channels, we started with the relative entropy which is a generalized divergence, and replaced states with channels and identity with the completely depolarising map and optimized overall input space of the channel. Similarly, we define the entropy of superchannels by taking the generalized divergence of channels as the starting point and replacing channels with superchannels and replacing the completely depolarising map with the supermap that serves its generalization in the space of supermaps, and then optimizing over all input space of the superchannel. Recall that the completely depolarizing map was defined as $\mathcal{R}\left(\mathsf{X}\right)=\operatorname{tr}(\mathsf{X})\mathbbm{1}$, where $\mathsf{X}$ was an arbitrary operator belonging to the input space. We now generalize this definition of the completely depolarizing map for the space of supermaps as follows. Let us denote by $\mathcal{R}^{(2)}$ a completely depolarising supermap.

### v2

@@ IV Sufficiency for the recoverability of quantum channels
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}^{2}_{1}$ is the fidelity of states $\rho$ and $\sigma$, $\mathsf{C}_{\mathcal{N}}^{\Psi}$ is the Choi state for channel $\mathcal{N}_{A\rightarrow A^{\prime}}$ with respect to $\Psi$ and $\mathcal{P}^{\mathrm{R}}$ is the recovery map defined by Eq. (70). $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{L}(C\otimes C^{\prime})$ is chosen in such a way that $\mathfrak{T}^{\prime}$ is a CPTP map (see Appendix B.4) and $\mathfrak{T}:\mathcal{L}\left(A\otimes A^{\prime}\right)\rightarrow\mathcal{L}\left(C\otimes C^{\prime}\right)$ is the representing map for superchannel $\Theta$ (see Remark 3 and Eq. (29)).
The detailed proof of this theorem is provided in Appendix  D while we sketch the proof of this theorem in the following. The basic idea behind the proof is the fact that a supermap induces a linear map, called the representing map, at the level of Choi matrices. The representing map $\mathfrak{T}$ of the superchannel $\Theta$ is in general a CP map that becomes trace preserving if the adjoint of $\Theta$ satisfies a simple condition (see Eq. (118) of Appendix III). This, in turn, implies that $\Theta^{*}$ is a completely depolarising map preserving (see Lemma 6) if $\Psi$ and $\Phi$ are maximally entangled states. Even if this trace preserving condition of the representing map $\mathfrak{T}$ is not satisfied, since we have $\Theta\in\mathsf{SCT}$, we can always replace the representing map $\mathfrak{T}$ with the following CPTP map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$ for all $\mathsf{X}\in\mathcal{L}(A\otimes A^{\prime})$. Here $\sigma_{0}\in\mathcal{L}(C\otimes C^{\prime})$ is chosen such that $\mathfrak{T}^{\prime}$ is a CPTP map (see Appendix B.4) and we further have $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$ and $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{M}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{M}})$. The recovery map $\mathcal{P}^{\mathrm{R}}$ is defined in such a way that it always recovers $\mathsf{C}^{\Psi}_{\mathcal{M}}$ from $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{M}})$.
@@
$\Theta^{\mathrm{R}}\circ\Theta\left(\mathcal{N}\right)=\mathcal{N}~{}~{}\text{and}~{}~{}\Theta^{\mathrm{R}}\circ\Theta\left(\mathcal{M}\right)=\mathcal{M}.$ (73)
@@
Let $\mathcal{N},\mathcal{M}\in\mathcal{L}\left(A,A^{\prime}\right)$ be two tele-covariant quantum channels (with respect to the same unitary representation of a group) (DW19a, , Definition 4), and let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a tele-covariance preserving superchannel and $\Theta\in\mathsf{SCT}$. Then we have
@@
Let $\mathcal{N}\in\mathcal{L}\left(A,A^{\prime}\right)$ be a tele-covariant channel, $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a tele-covariant channel preserving superchannel and $\Theta\in\mathsf{SCT}$. Then change in the entropy of quantum channel $\mathcal{N}$ is bounded as
@@
where $\mathsf{C}_{\mathcal{N}}$ is the Choi state for channel $\mathcal{N}$ and $\mathsf{T}\in\mathcal{L}\left(A\otimes A^{\prime},C\otimes C^{\prime}\right)$ is the representing map for superchannel $\Theta$ (see Remark 3 and Eq. (29)). $\mathsf{T}^{\prime}(\mathsf{X}):=\mathsf{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{L}(C\otimes C^{\prime})$ is chosen such that $\mathsf{T}^{\prime}$ is a CPTP map (see Appendix B.4). The recovery map $\widetilde{\mathcal{P}}^{\mathrm{R}}\left(\mathsf{X}\right)$ (see (BDW, 16, Eq. (36))) is given by $\widetilde{\mathcal{P}}^{\mathrm{R}}\left(\mathsf{X}\right)={\mathsf{T}^{\prime}}^{*}(\mathsf{X})+\operatorname{tr}\left(\left(\mathrm{id}-{\mathsf{T}^{\prime}}^{*}\right)\mathsf{X}\right)\xi$, where $\xi\in\mathcal{L}(C\otimes C^{\prime})$, $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$.
@@
Since $\mathcal{N}$ is a tele-covariant channel and $\Theta$ is a tele-covariant channel preserving superchannel, we have $S\left[\mathcal{N}\right]=S(\mathsf{C}_{\mathcal{N}})-\log\left(|A|\right)$ and $S\left[\Theta\left(\mathcal{N}\right)\right]=S\left(\mathsf{C}_{\Theta(\mathcal{N})}\right)-\log\left(|C|\right)$, where $\mathsf{C}_{\mathcal{N}}$ and $\mathsf{C}_{\Theta(\mathcal{N})}$ are the Choi states of $\mathcal{N}$ and $\Theta(\mathcal{N})$, respectively. Then the entropy difference becomes $S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]=S\left(\mathsf{C}_{\Theta(\mathcal{N})}\right)-S(\mathsf{C}_{\mathcal{N}})+\log\widetilde{d}$, where $\widetilde{d}:=|A|/|C|$. Also, $\mathsf{C}_{\mathcal{N}}$ and $\mathsf{C}_{\Theta(\mathcal{N})}$ are connected by the representing map $\mathsf{T}$ which is CP but not trace preserving, in general. So we replace it with a channel $\mathsf{T}^{\prime}$ such that $\mathsf{T}{{}^{\prime}}(\mathsf{C}_{\Psi})=\mathsf{T}(\mathsf{C}_{\Psi})$ holds when $|A|=|C|$ (see Eq. (20)). The entropy difference then is lower bounded by (see (BDW, 16, Eq. (37)), also DKSW (18))
@@ V Generalized divergence and entropy of superchannels
In Section III, we generalized the definition of entropy to quantum channels, where we argued that similar to the entropy of quantum states, the entropy of quantum channels can also be defined in terms of relative entropy between the given quantum channel and the completely depolarising map (see Definition 46). Here, we first extend the definition of generalized divergences to superchannels using the generalized divergence between two quantum channels. Recall that a map $\mathbf{D}:\mathcal{L}\left(A\right)_{+}\times\mathcal{L}\left(A\right)_{+}\rightarrow\mathbb{R}$ is called a generalized state divergence if it satisfies data processing inequality: $\mathbf{D}\left(\rho\;\middle\|\;\sigma\right)\geq\mathbf{D}\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)$, where $\rho,\sigma\in\mathcal{L}\left(A\right)_{+}$ and $\mathcal{N}:\mathcal{L}(A,A^{\prime})$ is a quantum channel. Given a generalized state divergence $\mathbf{D}$, we define the generalized divergence between a quantum channel $\mathcal{N}\in\mathcal{L}\left(A,A^{\prime}\right)$ and a CP map $\mathcal{M}\in\mathcal{L}\left(A,A^{\prime}\right)$ as
@@
where $R$ is a reference system of arbitrary dimension. Now, given a generalized divergence $\mathbf{D}$ between channels (cf. CMW (16); LKDW (18); Yua (19)), we define a divergence functional between a quantum superchannel $\Theta$ and a supermap $\Gamma$, denoted by $\mathbf{D}^{(2)}$, as follows.
@@
Given a generalized divergence $\mathbf{D}\left[\cdot\;\middle\|\;\cdot\right]$ between a quantum channel and a CP map (as in Eq. (77)), a divergence between a quantum superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ and a completely CP-preserving supermap $\Gamma:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is defined as
@@
where $\Lambda_{\text{pre}}\in\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(A\otimes R,B\otimes R\right)\right]$ and $\Lambda_{\text{post}}\in\mathscr{L}\left[\mathcal{L}\left(C\otimes R,D\otimes R\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ are preprocessing and postprocessing superchannels, respectively, similar to the preprocessing and postprocessing quantum channels of Eq. (13). $R$ is the reference system on which $\operatorname{id}$ acts. Looking at the mathematical similarity to quantum channels, we expect that the dilation equality, Eq. (79), will hold for all quantum super-superchannels. But, for the rest of this section, we restrict ourselves to the class $\mathsf{SSChan}$ of super-superchannels. For any pair of superchannels $\Theta,\Gamma\in\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]$, our definition of the divergence $\mathbf{D}^{(2)}$ for superchannels satisfies the monotonicity relation $\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\Gamma\right]\geq\mathbf{D}^{(2)}\left[\mathring{\mathbf{\Lambda}}\left(\Theta\right)\;\middle\|\;\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right]$ for all $\mathring{\mathbf{\Lambda}}\in\mathsf{SSChan}$. The proof goes as follows. For $\mathring{\mathbf{\Lambda}}\in\mathsf{SSChan}$, the superchannels $\mathring{\mathbf{\Lambda}}(\Theta)$ and $\mathring{\mathbf{\Lambda}}(\Gamma)$ can be written in terms of concatenation of superchannels given by Eq. (79). Using our definition of divergence between two superchannels, we can write
@@
In the following, we define a notion of generalized entropy for quantum superchannels based on the notion of the generalized divergence. Recall that generalized divergences $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ are functionals that take a pair of operators, say $A,B\geq 0$, as input and give a real number as output. One can define generalized entropies by taking $B$ to be the identity operator. For example, by taking $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ as relative entropy, and $B=\mathbbm{1}$, we get von Neumann entropy, which we have already discussed in the previous section. Also, by taking $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ as sandwiched Renyi divergence WWY (14); MLDS+ (13), and again $B=\mathbbm{1}$, one can get $\alpha$-Renyi entropy. To define the entropy of quantum channels, we started with the relative entropy which is a generalized divergence, and replaced states with channels and identity with the completely depolarising map and optimized overall input space of the channel. Similarly, we define the entropy of superchannels by taking the generalized divergence of channels as the starting point and replacing channels with superchannels and replacing the completely depolarising map with the supermap that serves its generalization in the space of supermaps, and then optimizing over all input space of the superchannel. Recall that the completely depolarizing map was defined as $\mathcal{R}\left(\mathsf{X}\right)=\operatorname{tr}(\mathsf{X})\mathbbm{1}$, where $\mathsf{X}$ was an arbitrary operator belonging to the input space. We now generalize this definition of the completely depolarizing map for the space of supermaps as follows. Let us denote by $\mathcal{R}^{(2)}$ a completely depolarising supermap.